OpenStudy (anonymous):

Forty-two is what percent of 70? A. 21% B. 60% C. 74% D. 167%

OpenStudy (saylilbaby):

B.60%

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

what r u people doing

OpenStudy (disco619):

\[\frac{ 42 }{ 70 } \times 100\] \[0.6 \times 100 = 60\] 60% is the answer, hope you get it. 

OpenStudy (mathstudent55):

We need to translate this word problem into a math equation and solve the equation. In math, "is" means an equal sign. "What" is an unknown, so we can use x. "Percent" means divide by 100. "Of" means times. Forty-two is what percent of 70? <--- English 42 = x /100 * 70 <--- Math You have this equation: \(42 = \dfrac{x}{100} \times 70\) \(42 = \dfrac{70x}{100} \) \(4200 = 70x\) \(70x = 4200\) \(x = 60\) Answer: 42 is 60% of 70.

OpenStudy (disco619):

Or that^ Haha